### v4

@@ Appendix E Summary of the CIQ Factors
The table provides summary of the estimated CIQ factors. In Panel A, we report descriptive statistics of the CIQ factors including their means, standard deviations, skewness, kurtosis and autocorrelation coefficients of order between one and three. In Panel B, we report correlations between CIQ factors. The data cover the period from January 1965 to December 2018.
| 0.1 | 0.15 | 0.2 | 0.3 | 0.4 | 0.5 | 0.6 | 0.7 | 0.8 | 0.85 | 0.9 | |
| Panel A: Descriptive statistics | |||||||||||
| Mean | -5.76 | -4.40 | -3.84 | -4.04 | -2.73 | 12.09 | -38.26 | -8.64 | -6.70 | -4.06 | -4.87 |
| St. Dev. | 0.15 | 0.16 | 0.18 | 0.25 | 0.47 | 1.43 | 1.33 | 0.35 | 0.22 | 0.19 | 0.17 |
| Skewness | -0.16 | -0.08 | -0.14 | 0.02 | 0.09 | 0.01 | -0.02 | -0.03 | -0.12 | -0.19 | 0.04 |
| Kurtosis | 4.38 | 4.43 | 4.79 | 5.75 | 5.86 | 5.51 | 5.56 | 7.02 | 7.17 | 7.32 | 6.72 |
| AR(1) | -0.41 | -0.40 | -0.44 | -0.46 | -0.48 | -0.36 | -0.34 | -0.48 | -0.43 | -0.41 | -0.38 |
| AR(2) | -0.10 | -0.08 | -0.05 | 0.01 | 0.04 | 0.05 | -0.05 | 0.03 | -0.05 | -0.07 | -0.11 |
| AR(3) | 0.15 | 0.11 | 0.06 | 0.01 | -0.03 | 0.01 | 0.00 | 0.02 | 0.06 | 0.07 | 0.07 |
| Panel B: Correlations | |||||||||||
| 0.1 | 1.00 | 0.96 | 0.93 | 0.84 | 0.66 | 0.09 | 0.15 | 0.15 | -0.09 | -0.20 | -0.29 |
| 0.15 | 1.00 | 0.97 | 0.90 | 0.75 | 0.15 | 0.23 | 0.24 | 0.00 | -0.11 | -0.22 | |
| 0.2 | 1.00 | 0.95 | 0.81 | 0.19 | 0.28 | 0.34 | 0.10 | -0.02 | -0.13 | ||
| 0.3 | 1.00 | 0.93 | 0.28 | 0.37 | 0.50 | 0.28 | 0.16 | 0.04 | |||
| 0.4 | 1.00 | 0.43 | 0.50 | 0.68 | 0.49 | 0.38 | 0.26 | ||||
| 0.5 | 1.00 | 0.72 | 0.49 | 0.43 | 0.39 | 0.34 | |||||
| 0.6 | 1.00 | 0.59 | 0.50 | 0.46 | 0.38 | ||||||
| 0.7 | 1.00 | 0.93 | 0.86 | 0.76 | |||||||
| 0.8 | 1.00 | 0.97 | 0.91 | ||||||||
| 0.85 | 1.00 | 0.95 | |||||||||
| 0.9 | 1.00 | ||||||||||
E.1 CIQ Shocks and the Real Economy
A large body of literature shows that firm-level shocks can be linked to individual households via channels such as wages, employment and wealth due to incomplete insurance (Harris and Holmstrom, 1982; Brown and Medoff, 1989; Berk et al., 2010; Heathcote et al., 2014; Herskovic et al., 2016). This suggests that common quantile factors (CIQ) in firm returns may proxy for household risk. We test this hypothesis by examining the correlations between the innovations in the CIQ and the cross-sectional outcomes of firm employment growth, earnings growth (by place of work), house price growth and firm sales growth. For each outcome, we compute dispersion (interquartile range, etc.) and key percentiles (20th, 40th, 60th and 80th) to capture heterogeneous results.
Table 13 shows the correlations between the changes in CIQ and the changes in these distributional measures. The strongest link is shown by household earnings growth: a lower-tail CIQ shock is strongly associated with declines in earnings growth, whereas an upper-tail shock coincides with higher earnings growth. Sales growth (Panel E) exhibits greater asymmetry: an upper-tail shock improves the performance of the weakest firms, thereby compressing the distribution, whereas a lower-tail shock hurts these firms, thereby widening the distribution. In contrast, a volatility shock (CIV) increases dispersion and depresses all sales growth percentiles. Employment growth (Panel A) and house prices (Panels C–D) exhibit weaker yet consistent patterns. A downside CIQ shock increases employment growth dispersion and exacerbates job losses at the lower end of the scale, whereas an upside shock mitigates the worst outcomes. House price growth only responds to downside CIQ shocks, indicating that housing wealth risk is mainly driven by common downturns. Overall, these results confirm that CIQ risk is associated with fluctuations in household earnings, wealth and sales. The effects differ across quantiles, indicating heterogeneous effects that would be missed by focusing only on volatility (CIV). This highlights the heterogeneous transmission of risk from firms to households and implies that asset pricing models should incorporate quantile-dependent risk (e.g. via quantile-based preferences, as discussed in de Castro and Galvao (2019)) rather than relying solely on variance.
This table presents the correlations between CIQ risks and changes in cross-sectional measures of employment (Panel A), earnings (Panel B), house-price (Panels C and D) growth, and sales growth (Panel E). It reports dispersion measures (inter-quartile range, max–min, standard deviation), selected percentiles , and the cross-sectional average.
| std | Avg. | |||||||
| Panel A: Annual employment growth | ||||||||
| Lower-tail CIQ | 0.46 ** | 0.33 | 0.27 | -0.41 * | -0.61 *** | 0.13 | 0.25 | -0.21 |
| Central CIQ | 0.18 | 0.30 | 0.34 | 0.13 | 0.64 *** | 0.35 | 0.49 ** | 0.38 * |
| Upper-tail CIQ | -0.38 * | -0.01 | -0.18 | 0.56 *** | 0.64 *** | 0.25 | 0.05 | 0.49 ** |
| CIV | -0.10 | -0.01 | -0.09 | -0.27 | -0.25 | -0.52 *** | -0.47 ** | -0.51 *** |
| Panel B: Annual earnings growth | ||||||||
| Lower-tail CIQ | -0.19 | -0.42 * | -0.37 | -0.38 * | -0.46 ** | -0.50 *** | -0.56 *** | -0.50 *** |
| Central CIQ | -0.11 | 0.28 | -0.04 | 0.41 * | 0.45 ** | 0.46 ** | 0.40 * | 0.41 * |
| Upper-tail CIQ | 0.12 | 0.49 ** | 0.20 | 0.45 ** | 0.57 *** | 0.62 *** | 0.59 *** | 0.49 ** |
| CIV | 0.02 | -0.0 | -0.15 | -0.48 ** | -0.54 *** | -0.54 *** | -0.47 ** | -0.51 *** |
| Panel C: Monthly house‐price growth (nine areas, seasonally adjusted) | ||||||||
| Lower-tail CIQ | -0.03 | 0.06 | 0.04 | 0.13 *** | 0.06 | 0.08 | 0.13 *** | 0.11 ** |
| Central CIQ | -0.02 | 0.09 | 0.07 | 0.05 | 0.05 | 0.00 | 0.04 | 0.05 |
| Upper-tail CIQ | -0.02 | 0.01 | 0.01 | -0.01 | 0.00 | -0.02 | -0.00 | 0.00 |
| CIV | 0.03 | -0.04 | -0.02 | -0.12 ** | -0.12 ** | -0.06 | -0.12 ** | -0.13 *** |
| Panel D: Quarterly house-price growth (100 largest MSAs, seasonally adjusted) | ||||||||
| Lower-tail CIQ | 0.04 | 0.15 | 0.13 | 0.16 * | 0.21 ** | 0.12 | 0.19 ** | 0.18 * |
| Central CIQ | -0.06 | 0.10 | 0.14 | 0.06 | 0.05 | 0.01 | 0.01 | 0.03 |
| Upper-tail CIQ | -0.04 | -0.07 | -0.07 | 0.00 | 0.05 | -0.03 | 0.01 | 0.05 |
| CIV | -0.08 | 0.12 | -0.01 | -0.11 | -0.10 | -0.08 | -0.14 | -0.12 |
| Panel E: Sales growth | ||||||||
| Lower-tail CIQ | -0.26 ** | 0.02 | -0.15 | 0.13 | 0.06 | -0.08 | -0.12 | 0.05 |
| Central CIQ | -0.02 | 0.05 | 0.04 | -0.00 | -0.02 | -0.04 | -0.01 | -0.04 |
[…]

### v5

@@ Appendix G Summary of the CIQ Factors
The table provides summary of the estimated CIQ factors. In Panel A, we report descriptive statistics of the CIQ factors including their means, standard deviations, skewness, kurtosis and autocorrelation coefficients of order between one and three. In Panel B, we report correlations between CIQ factors. The data cover the period from January 1968 to December 2024.
| 0.05 | 0.1 | 0.15 | 0.2 | 0.25 | 0.3 | 0.35 | 0.4 | 0.45 | 0.5 | 0.55 | 0.6 | 0.65 | 0.7 | 0.75 | 0.8 | 0.85 | 0.9 | 0.95 | |
| Panel A: Descriptive statistics | |||||||||||||||||||
| Mean | -5.83 | -5.17 | -3.45 | -3.21 | -4.02 | -4.07 | -6.18 | -9.23 | -35.29 | -41.48 | -78.70 | -52.38 | -16.06 | -9.55 | -7.92 | -8.34 | -5.93 | -6.25 | 0.57 |
| St. Dev. | 0.15 | 0.16 | 0.17 | 0.19 | 0.22 | 0.26 | 0.33 | 0.49 | 1.27 | 1.46 | 1.46 | 1.32 | 0.57 | 0.36 | 0.28 | 0.23 | 0.20 | 0.18 | 0.17 |
| Skewness | -0.10 | -0.06 | 0.06 | 0.06 | 0.22 | 0.15 | 0.25 | 0.27 | -0.27 | -0.35 | -0.30 | -0.06 | 0.19 | 0.19 | 0.03 | -0.05 | -0.19 | -0.01 | 0.26 |
| Kurtosis | 4.42 | 4.21 | 4.57 | 5.09 | 5.80 | 5.69 | 6.06 | 6.09 | 8.09 | 6.41 | 6.55 | 5.46 | 6.90 | 7.20 | 6.63 | 6.86 | 6.78 | 6.18 | 5.57 |
| AR(1) | -0.36 | -0.40 | -0.41 | -0.44 | -0.43 | -0.46 | -0.46 | -0.46 | -0.29 | -0.27 | -0.27 | -0.33 | -0.47 | -0.46 | -0.44 | -0.42 | -0.40 | -0.37 | -0.33 |
| AR(2) | -0.14 | -0.08 | -0.05 | -0.02 | -0.00 | 0.03 | 0.03 | 0.05 | 0.07 | 0.07 | 0.02 | -0.01 | 0.01 | -0.00 | -0.05 | -0.07 | -0.10 | -0.13 | -0.13 |
| AR(3) | 0.17 | 0.12 | 0.08 | 0.03 | 0.00 | -0.02 | -0.02 | -0.04 | -0.07 | -0.02 | -0.00 | -0.01 | -0.00 | 0.02 | 0.06 | 0.07 | 0.09 | 0.09 | 0.08 |
| Panel B: Correlations | |||||||||||||||||||
| 0.05 | 1.00 | 0.93 | 0.89 | 0.85 | 0.79 | 0.72 | 0.65 | 0.54 | 0.19 | 0.07 | 0.05 | 0.11 | 0.15 | 0.05 | -0.06 | -0.15 | -0.24 | -0.33 | -0.39 |
| 0.10 | 1.00 | 0.97 | 0.93 | 0.89 | 0.84 | 0.78 | 0.67 | 0.28 | 0.14 | 0.11 | 0.18 | 0.25 | 0.16 | 0.03 | -0.06 | -0.17 | -0.26 | -0.34 | |
| 0.15 | 1.00 | 0.98 | 0.95 | 0.91 | 0.85 | 0.76 | 0.35 | 0.19 | 0.16 | 0.26 | 0.35 | 0.26 | 0.13 | 0.03 | -0.08 | -0.19 | -0.27 | ||
| 0.20 | 1.00 | 0.98 | 0.95 | 0.91 | 0.82 | 0.40 | 0.22 | 0.20 | 0.32 | 0.44 | 0.36 | 0.23 | 0.13 | 0.01 | -0.10 | -0.20 | |||
| 0.25 | 1.00 | 0.98 | 0.95 | 0.87 | 0.45 | 0.26 | 0.23 | 0.36 | 0.51 | 0.43 | 0.31 | 0.21 | 0.09 | -0.03 | -0.14 | ||||
| 0.30 | 1.00 | 0.98 | 0.92 | 0.51 | 0.31 | 0.28 | 0.41 | 0.58 | 0.52 | 0.40 | 0.31 | 0.19 | 0.07 | -0.07 | |||||
| 0.35 | 1.00 | 0.96 | 0.58 | 0.38 | 0.34 | 0.47 | 0.66 | 0.61 | 0.50 | 0.41 | 0.29 | 0.16 | 0.02 | ||||||
| 0.40 | 1.00 | 0.69 | 0.47 | 0.43 | 0.55 | 0.74 | 0.69 | 0.60 | 0.52 | 0.41 | 0.29 | 0.14 | |||||||
| 0.45 | 1.00 | 0.77 | 0.72 | 0.71 | 0.67 | 0.61 | 0.56 | 0.51 | 0.45 | 0.36 | 0.26 | ||||||||
| 0.50 | 1.00 | 0.90 | 0.73 | 0.53 | 0.47 | 0.44 | 0.41 | 0.38 | 0.31 | 0.24 | |||||||||
| 0.55 | 1.00 | 0.80 | 0.52 | 0.47 | 0.44 | 0.41 | 0.38 | 0.32 | 0.26 | ||||||||||
| 0.60 | 1.00 | 0.69 | 0.61 | 0.57 | 0.53 | 0.48 | 0.40 | 0.29 | |||||||||||
| 0.65 | 1.00 | 0.96 | 0.92 | 0.87 | 0.80 | 0.69 | 0.54 | ||||||||||||
| 0.70 | 1.00 | 0.97 | 0.93 | 0.87 | 0.77 | 0.62 | |||||||||||||
| 0.75 | 1.00 | 0.97 | 0.93 | 0.86 | 0.72 | ||||||||||||||
| 0.80 | 1.00 | 0.97 | 0.91 | 0.78 | |||||||||||||||
| 0.85 | 1.00 | 0.96 | 0.85 | ||||||||||||||||
| 0.90 | 1.00 | 0.91 | |||||||||||||||||
| 0.95 | 1.00 | ||||||||||||||||||
Appendix H Predicting the CIQ Premium
The table reports predictive regressions of the high-minus-low equal-weighted portfolio returns from decile sorts on the exposures to the factor. We regress the portfolio’s one-month-ahead returns on the lagged market return (), on the CBOE variance factor () and on its two components: the conditional variance of stock returns (CV) and the equity variance premium (VP) based on the procedure of Bekaert and Hoerova (2014). Coefficients are scaled to capture the effect of one-standard-deviation increase in the factor on the annualized portfolio return in percentage points. -statistics (in parentheses) are based on standard errors of Newey and West (1994). The data cover the period between February 1990 and January 2022.
| (1) | (2) | (3) | (4) | (5) | (6) | |
| 6.10 | ||||||
| 7.28 | 8.61 | 7.81 | ||||
| (2.27) | (2.71) | (2.34) | ||||
| 4.30 | -1.86 | 1.23 | ||||
| (1.37) | (-0.84) | (0.43) | ||||
| 2.74 | 5.36 | |||||
| (1.14) | (2.23) | |||||
| IS | 2.22 | 3.16 | 1.10 | 0.45 | 3.26 | 4.58 |
| OOS | 0.90 | 2.44 | -0.77 | -2.45 | -0.10 | -0.83 |
| OOS CT | 1.49 | 2.44 | 0.42 | -0.97 | 2.34 | 2.00 |
Appendix I CIQ premia
The table contains annualized out-of-sample excess returns of ten portfolios sorted on the exposure to the CIQ() factors. We also report returns of the high minus low (H - L) portfolios, their -statistics, and annualized alphas with respect to the six-factor model of Fama and French (2018). All -statistics are based on the correction of Newey and West (1994). The data cover the period from January 1968 to December 2024. Each month, we use all the CRSP stocks with at least 48 monthly observations over the last 60 months, and exclude penny stocks with prices below $1.
| Low | 2 | 3 | 4 | 5 | 6 | 7 | 8 | 9 | High | H - L | -stat | -stat | ||
| Equal-weighted | ||||||||||||||
| 0.05 | 5.68 | 8.17 | 9.25 | 9.51 | 9.93 | 9.55 | 10.08 | 10.95 | 11.55 | 10.87 | 5.19 | 2.97 | 5.35 | 3.05 |
| 0.10 | 5.14 | 7.91 | 8.76 | 9.59 | 9.50 | 10.28 | 10.68 | 10.72 | 11.98 | 10.96 | 5.82 | 3.29 | 6.07 | 3.68 |
| 0.15 | 4.78 | 8.02 | 9.13 | 9.79 | 10.01 | 9.69 | 10.60 | 10.72 | 11.22 | 11.58 | 6.80 | 3.76 | 7.13 | 4.25 |
| 0.20 | 4.36 | 7.94 | 9.47 | 9.63 | 9.46 | 10.36 | 10.41 | 10.74 | 11.39 | 11.77 | 7.41 | 4.30 | 8.31 | 4.97 |
| 0.25 | 4.78 | 7.76 | 9.26 | 9.93 | 10.20 | 10.15 | 10.41 | 10.45 | 11.06 | 11.54 | 6.76 | 4.09 | 7.86 | 5.10 |
| 0.30 | 4.86 | 7.98 | 9.14 | 9.89 | 10.03 | 10.51 | 10.35 | 10.87 | 10.64 | 11.27 | 6.41 | 3.90 | 7.86 | 5.05 |
| 0.35 | 4.93 | 8.37 | 9.63 | 9.65 | 9.74 | 10.47 | 10.46 | 10.40 | 11.00 | 10.89 | 5.96 | 3.83 | 7.46 | 4.81 |
| 0.40 | 5.06 | 8.57 | 9.49 | 9.56 | 9.73 | 10.28 | 10.77 | 10.84 | 10.72 | 10.53 | 5.46 | 3.72 | 6.79 | 4.60 |
| 0.45 | 5.80 | 7.95 | 10.31 | 9.34 | 10.00 | 10.28 | 10.49 | 10.46 | 10.71 | 10.19 | 4.38 | 2.52 | 4.80 | 2.60 |
| 0.50 | 6.90 | 9.05 | 10.35 | 9.05 | 10.23 | 9.85 | 10.31 | 10.24 | 10.30 | 9.26 | 2.36 | 1.36 | 1.68 | 0.85 |
| 0.55 | 6.36 | 8.85 | 10.03 | 9.27 | 10.18 | 9.99 | 10.43 | 10.28 | 10.53 | 9.61 | 3.25 | 1.86 | 2.27 | 1.09 |
| 0.60 | 7.16 | 9.26 | 10.92 | 9.61 | 10.32 | 10.00 | 9.93 | 9.58 | 9.95 | 8.82 | 1.66 | 1.01 | 0.95 | 0.51 |
| 0.65 | 6.84 | 9.89 | 9.69 | 10.31 | 9.77 | 9.62 | 9.97 | 10.08 | 10.24 | 9.13 | 2.29 | 1.51 | 1.73 | 1.08 |
| 0.70 | 7.80 | 9.37 | 10.39 | 10.26 | 9.49 | 9.72 | 10.05 | 9.45 | 10.21 | 8.79 | 0.99 | 0.63 | -0.01 | -0.01 |
| 0.75 | 8.26 | 9.79 | 10.35 | 9.78 | 9.81 | 9.88 | 9.81 | 9.78 | 9.04 | 9.03 | 0.77 | 0.50 | -0.49 | -0.30 |
| 0.80 | 8.86 | 10.09 | 10.46 | 9.62 | 9.73 | 9.70 | 9.90 | 9.00 | 9.74 | 8.43 | -0.42 | -0.26 | -1.60 | -0.92 |
| 0.85 | 9.07 | 10.28 | 9.79 | 10.03 | 9.67 | 9.59 | 9.77 | 9.69 | 9.02 | 8.63 | -0.44 | -0.26 | -1.65 | -0.90 |
| 0.90 | 9.31 | 10.25 | 9.88 | 9.97 | 9.33 | 10.22 | 10.04 | 9.34 | 8.76 | 8.43 | -0.89 | -0.50 | -2.16 | -1.18 |
| 0.95 | 9.65 | 10.44 | 10.28 | 9.67 | 10.03 | 9.36 | 10.41 | 8.78 | 9.14 | 7.77 | -1.88 | -0.97 | -3.20 | -1.75 |
| Value-weighted | ||||||||||||||
| 0.05 | 6.39 | 5.95 | 6.45 | 7.62 | 6.64 | 7.58 | 7.52 | 9.51 | 9.96 | 8.75 | 2.36 | 0.96 | 0.70 | 0.28 |
| 0.10 | 5.16 | 6.12 | 5.98 | 7.81 | 6.80 | 7.95 | 7.83 | 8.66 | 10.35 | 8.96 | 3.80 | 1.51 | 3.02 | 1.22 |
| 0.15 | 4.41 | 6.59 | 7.11 | 6.79 | 7.47 | 7.75 | 8.84 | 7.49 | 10.26 | 9.76 | 5.35 | 2.05 | 5.21 | 2.03 |
| 0.20 | 3.35 | 7.29 | 6.65 | 6.56 | 7.17 | 8.11 | 7.98 | 8.24 | 9.81 | 10.73 | 7.37 | 2.73 | 7.86 | 2.84 |
| 0.25 | 4.40 | 6.68 | 6.80 | 6.88 | 8.15 | 7.84 | 7.90 | 7.30 | 9.47 | 10.35 | 5.95 | 2.27 | 6.57 | 2.43 |
| 0.30 | 4.05 | 7.65 | 6.98 | 7.36 | 7.71 | 8.07 | 7.51 | 7.06 | 8.73 | 10.23 | 6.18 | 2.27 | 6.69 | 2.37 |
| 0.35 | 5.18 | 7.16 | 7.45 | 7.36 | 7.22 | 8.19 | 7.37 | 6.77 | 8.42 | 9.86 | 4.68 | 1.77 | 5.83 | 2.18 |
| 0.40 | 4.81 | 7.57 | 6.50 | 8.16 | 7.56 | 7.95 | 6.73 | 7.37 | 7.47 | 9.99 | 5.18 | 2.15 | 6.96 | 2.83 |
| 0.45 | 4.43 | 5.41 | 7.10 | 6.92 | 7.76 | 8.16 | 8.26 | 7.29 | 7.58 | 10.13 | 5.70 | 2.44 | 6.55 | 2.71 |
| 0.50 | 5.02 | 5.18 | 6.63 | 6.58 | 7.51 | 8.43 | 8.15 | 7.73 | 6.49 | 9.38 | 4.36 | 1.81 | 3.99 | 1.56 |
| 0.55 | 5.03 | 4.72 | 6.90 | 6.21 | 7.85 | 7.60 | 8.31 | 8.53 | 6.86 | 9.06 | 4.03 | 1.60 | 3.46 | 1.22 |
| 0.60 | 4.61 | 6.73 | 7.90 | 6.50 | 7.43 | 7.47 | 8.07 | 7.50 | 6.94 | 7.90 | 3.29 | 1.40 | 2.45 | 1.02 |
| 0.65 | 4.62 | 7.98 | 7.98 | 8.94 | 7.32 | 7.17 | 6.40 | 7.22 | 6.92 | 7.32 | 2.70 | 1.22 | 2.30 | 0.95 |
| 0.70 | 5.64 | 7.48 | 8.84 | 8.98 | 7.55 | 7.34 | 6.93 | 6.89 | 6.49 | 7.08 | 1.44 | 0.59 | 0.87 | 0.36 |
| 0.75 | 4.86 | 8.91 | 7.78 | 8.74 | 7.98 | 7.58 | 6.48 | 6.63 | 5.38 | 6.69 | 1.83 | 0.76 | 0.90 | 0.37 |
| 0.80 | 7.23 | 9.11 | 8.15 | 7.07 | 8.11 | 7.84 | 6.81 | 6.37 | 5.87 | 6.33 | -0.90 | -0.36 | 0.03 | 0.01 |
| 0.85 | 6.91 | 8.92 | 7.71 | 6.76 | 8.25 | 7.22 | 7.36 | 7.01 | 6.01 | 5.63 | -1.27 | -0.51 | 0.08 | 0.03 |
| 0.90 | 6.74 | 8.84 | 7.44 | 6.75 | 7.98 | 7.80 | 6.82 | 6.69 | 6.79 | 5.49 | -1.24 | -0.50 | 0.03 | 0.01 |
| 0.95 | 5.99 | 8.62 | 7.72 | 8.08 | 7.15 | 7.39 | 7.68 | 6.16 | 6.03 | 5.14 | -0.85 | -0.31 | 0.09 | 0.04 |
The table contains annualized out-of-sample excess returns of five portfolios sorted on the exposure to the CIQ() factors. We also report returns of the high minus low (H - L) portfolios, their -statistics, and annualized alphas with respect to the six-factor model of Fama and French (2018). All -statistics are based on the correction of Newey and West (1994). The data cover the period from January 1968 to December 2024. Each month, we use all the CRSP stocks with at least 48 monthly observations over the last 60 months, and exclude penny stocks with prices below $1.
| Low | 2 | 3 | 4 | High | H - L | -stat | -stat | ||
| Equal-weighted | |||||||||
| 0.05 | 6.93 | 9.38 | 9.74 | 10.51 | 11.21 | 4.28 | 2.93 | 4.91 | 3.36 |
| 0.10 | 6.53 | 9.17 | 9.89 | 10.70 | 11.47 | 4.95 | 3.28 | 5.66 | 4.09 |
| 0.15 | 6.40 | 9.46 | 9.85 | 10.66 | 11.40 | 5.00 | 3.33 | 5.89 | 4.33 |
| 0.20 | 6.15 | 9.55 | 9.91 | 10.57 | 11.58 | 5.43 | 3.76 | 6.83 | 5.11 |
| 0.25 | 6.27 | 9.59 | 10.18 | 10.43 | 11.30 | 5.03 | 3.67 | 6.65 | 5.50 |
| 0.30 | 6.42 | 9.51 | 10.27 | 10.61 | 10.96 | 4.54 | 3.43 | 6.23 | 5.25 |
| 0.35 | 6.65 | 9.64 | 10.11 | 10.43 | 10.94 | 4.29 | 3.49 | 6.07 | 5.43 |
| 0.40 | 6.81 | 9.52 | 10.01 | 10.80 | 10.62 | 3.81 | 3.33 | 5.42 | 5.07 |
| 0.45 | 6.88 | 9.82 | 10.14 | 10.48 | 10.45 | 3.57 | 2.72 | 4.18 | 3.06 |
| 0.50 | 7.97 | 9.70 | 10.04 | 10.27 | 9.78 | 1.81 | 1.36 | 1.56 | 1.04 |
| 0.55 | 7.60 | 9.65 | 10.09 | 10.36 | 10.07 | 2.47 | 1.87 | 1.94 | 1.25 |
| 0.60 | 8.21 | 10.26 | 10.16 | 9.76 | 9.38 | 1.17 | 0.98 | 0.87 | 0.64 |
| 0.65 | 8.36 | 10.00 | 9.70 | 10.02 | 9.68 | 1.32 | 1.12 | 1.18 | 0.96 |
| 0.70 | 8.59 | 10.32 | 9.61 | 9.75 | 9.50 | 0.92 | 0.75 | 0.26 | 0.20 |
| 0.75 | 9.03 | 10.07 | 9.85 | 9.79 | 9.03 | 0.01 | 0.01 | -0.85 | -0.64 |
| 0.80 | 9.48 | 10.03 | 9.72 | 9.45 | 9.09 | -0.39 | -0.29 | -1.36 | -0.99 |
| 0.85 | 9.67 | 9.91 | 9.63 | 9.73 | 8.82 | -0.85 | -0.60 | -1.91 | -1.33 |
| 0.90 | 9.78 | 9.93 | 9.77 | 9.69 | 8.59 | -1.19 | -0.79 | -2.31 | -1.59 |
| 0.95 | 10.05 | 9.98 | 9.69 | 9.60 | 8.45 | -1.59 | -0.98 | -2.90 | -2.03 |
| Value-weighted | |||||||||
| 0.05 | 6.03 | 7.01 | 6.93 | 8.38 | 9.46 | 3.42 | 1.65 | 2.43 | 1.28 |
| 0.10 | 5.77 | 6.96 | 7.40 | 8.11 | 9.94 | 4.17 | 2.04 | 3.75 | 1.96 |
| 0.15 | 5.91 | 6.91 | 7.62 | 8.21 | 10.02 | 4.12 | 2.00 | 4.28 | 2.31 |
| 0.20 | 5.90 | 6.61 | 7.69 | 7.98 | 9.96 | 4.06 | 2.02 | 4.82 | 2.71 |
| 0.25 | 6.01 | 6.92 | 8.02 | 7.32 | 9.61 | 3.60 | 1.84 | 4.49 | 2.51 |
| 0.30 | 6.37 | 6.99 | 8.02 | 7.11 | 9.13 | 2.77 | 1.45 | 3.72 | 2.08 |
| 0.35 | 6.48 | 7.28 | 7.63 | 6.95 | 8.87 | 2.38 | 1.27 | 3.84 | 2.19 |
| 0.40 | 6.52 | 7.29 | 7.55 | 6.98 | 8.24 | 1.72 | 0.95 | 3.25 | 1.83 |
| 0.45 | 4.92 | 6.88 | 7.86 | 7.64 | 8.82 | 3.90 | 2.31 | 4.96 | 2.72 |
| 0.50 | 5.06 | 6.55 | 7.89 | 8.04 | 7.57 | 2.51 | 1.46 | 2.25 | 1.13 |
| 0.55 | 4.73 | 6.46 | 7.65 | 8.40 | 7.62 | 2.89 | 1.70 | 2.75 | 1.40 |
| 0.60 | 6.07 | 6.99 | 7.43 | 7.66 | 7.23 | 1.16 | 0.70 | 0.74 | 0.40 |
| 0.65 | 6.80 | 8.40 | 7.13 | 6.61 | 7.21 | 0.42 | 0.25 | 0.66 | 0.36 |
| 0.70 | 6.97 | 8.95 | 7.31 | 6.85 | 6.70 | -0.27 | -0.16 | -0.64 | -0.36 |
| 0.75 | 7.78 | 8.44 | 7.61 | 6.53 | 6.04 | -1.74 | -1.04 | -2.24 | -1.24 |
| 0.80 | 8.34 | 7.75 | 7.74 | 6.61 | 6.34 | -2.00 | -1.12 | -2.15 | -1.13 |
| 0.85 | 8.53 | 7.31 | 7.63 | 7.21 | 6.07 | -2.47 | -1.31 | -2.29 | -1.20 |
| 0.90 | 8.24 | 6.88 | 8.02 | 6.76 | 6.51 | -1.73 | -0.88 | -1.11 | -0.59 |
| 0.95 | 7.83 | 8.15 | 7.30 | 6.98 | 5.82 | -2.01 | -0.94 | -1.46 | -0.76 |
The table provides time-series correlations between the CIQ premia. We use the high-minus-low equal-weighted portfolios from decile sorts based on the exposures to the CIQ factors and report the correlations for range of values of . The data cover the period from January 1968 to December 2024.
| 0.05 | 0.1 | 0.15 | 0.2 | 0.25 | 0.3 | 0.35 | 0.4 | 0.45 | 0.5 | 0.55 | 0.6 | 0.65 | 0.7 | 0.75 | 0.8 | 0.85 | 0.9 | 0.95 | |
| 0.05 | 1.00 | 0.93 | 0.90 | 0.87 | 0.82 | 0.77 | 0.69 | 0.56 | 0.25 | 0.15 | 0.16 | 0.14 | 0.04 | -0.16 | -0.32 | -0.46 | -0.55 | -0.60 | -0.66 |
| 0.10 | 1.00 | 0.96 | 0.93 | 0.90 | 0.85 | 0.78 | 0.67 | 0.35 | 0.22 | 0.24 | 0.23 | 0.14 | -0.06 | -0.23 | -0.37 | -0.47 | -0.54 | -0.62 | |
| 0.15 | 1.00 | 0.96 | 0.93 | 0.89 | 0.84 | 0.72 | 0.41 | 0.28 | 0.30 | 0.26 | 0.18 | -0.02 | -0.19 | -0.33 | -0.43 | -0.51 | -0.61 | ||
| 0.20 | 1.00 | 0.97 | 0.94 | 0.88 | 0.79 | 0.44 | 0.30 | 0.32 | 0.32 | 0.26 | 0.06 | -0.12 | -0.27 | -0.38 | -0.47 | -0.58 | |||
| 0.25 | 1.00 | 0.97 | 0.93 | 0.84 | 0.49 | 0.33 | 0.36 | 0.36 | 0.34 | 0.13 | -0.04 | -0.20 | -0.31 | -0.41 | -0.53 | ||||
| 0.30 | 1.00 | 0.96 | 0.89 | 0.53 | 0.36 | 0.39 | 0.43 | 0.42 | 0.23 | 0.05 | -0.11 | -0.23 | -0.34 | -0.47 | |||||
| 0.35 | 1.00 | 0.94 | 0.59 | 0.41 | 0.44 | 0.47 | 0.50 | 0.33 | 0.17 | 0.02 | -0.10 | -0.21 | -0.35 | ||||||
| 0.40 | 1.00 | 0.71 | 0.51 | 0.53 | 0.57 | 0.60 | 0.46 | 0.32 | 0.18 | 0.07 | -0.04 | -0.19 | |||||||
| 0.45 | 1.00 | 0.77 | 0.77 | 0.69 | 0.58 | 0.47 | 0.42 | 0.34 | 0.28 | 0.20 | 0.09 | ||||||||
| 0.50 | 1.00 | 0.91 | 0.72 | 0.55 | 0.48 | 0.43 | 0.36 | 0.33 | 0.26 | 0.16 | |||||||||
| 0.55 | 1.00 | 0.75 | 0.59 | 0.51 | 0.45 | 0.39 | 0.35 | 0.28 | 0.16 | ||||||||||
| 0.60 | 1.00 | 0.75 | 0.65 | 0.56 | 0.49 | 0.42 | 0.33 | 0.17 | |||||||||||
| 0.65 | 1.00 | 0.92 | 0.82 | 0.73 | 0.62 | 0.50 | 0.32 | ||||||||||||
| 0.70 | 1.00 | 0.94 | 0.87 | 0.78 | 0.68 | 0.51 | |||||||||||||
| 0.75 | 1.00 | 0.95 | 0.89 | 0.81 | 0.68 | ||||||||||||||
| 0.80 | 1.00 | 0.96 | 0.90 | 0.79 | |||||||||||||||
| 0.85 | 1.00 | 0.95 | 0.86 | ||||||||||||||||
| 0.90 | 1.00 | 0.92 | |||||||||||||||||
| 0.95 | 1.00 |
Appendix J CIQ Betas and Firm Characteristics
This table reports Fama–MacBeth cross-sectional regressions explaining firm-level exposure to innovations in the upper-tail CIQ factor. The dependent variable is each stock’s rolling 60-month beta with respect to . Firm characteristics are ranked cross-sectionally each month and linearly scaled to lie between and . Regressions are estimated monthly, and reported coefficients are time-series averages. -statistics based on Newey and West (1994) are reported in parentheses. Firm-level characteristics are from Freyberger et al. (2020) and Kim et al. (2020). The sample covers the period from January 1968 to December 2018.
| (1) | (2) | (3) | (4) | (5) | |
| Liquidity | Funding | Risk | Vulnerability | Full | |
| Liquidity fragility | |||||
| dto | 0.01 | 0.00 | |||
| (2.35) | (1.75) | ||||
| lturnover | 0.10 | 0.08 | |||
[…]